# MorphAny3D: Unleashing the Power of Structured Latent in 3D Morphing

### v1

@@ 1 Introduction
Morphing [2Dsurvey1, 2Dsurvey2, 2D3Dsurvey] refers to the visual effect of seamlessly transforming a source into a target through a smooth, plausible, and aesthetic deformation sequence. This foundational technique is widely used in animation, film, and game design to enhance creative expression. Based on input dimensionality, morphing is categorized into 2D morphing [2Dmorphing1, 2Ddatamorphing1, diffmorpher, freemorph] and 3D morphing [lmkmap1, datamap1, morphflow, gaussianmorphing, 3Dmorpher]. While 2D morphing has advanced significantly with diffusion models [diffusion_ddim, diffusion_ddpm, diffusion_sde, stable_diffusion], 3D morphing remains challenging due to the inherent complexity of modeling smooth and reasonable deformations in three dimensions.
@@
Most existing 3D morphing approaches follow a two-stage paradigm: (1) establishing dense correspondences [3Dmatch_survey1, 3Dmatch_survey2] between source and target objects using hand-crafted sparse landmarks [lmkmap1, lmkmap2], functional maps [functionmap3], optimal transport [opt_trans1], data priors [datamap1], or extracted 2D features [featuremap2]; and (2) interpolating between these correspondences to generate intermediate shapes. While effective in constrained settings, such matching-based methods face several critical limitations. First, they typically prioritize geometric deformation while ignoring the concurrent evolution of textures. Second, they often exhibit limited generalization, particularly in cross-category transformations (e.g., morphing a chair into a car). As shown in Fig. 2-(a), inaccurate correspondence estimation frequently leads to structurally implausible morphing results.
@@
Recent advances in feed-forward 3D generation frameworks [lgm, wonder3d, 3DShape2VecSet, hunyuan3d_v2, clay, trellis, direct3d] have substantially improved the quality and fidelity of 3D synthesis. Notably, Trellis [trellis] achieves diverse and high-fidelity 3D generation through its Structured Latent (SLAT) representation. Owing to SLAT’s explicit and regular structure, it shows strong potential for training-free applications such as 3D editing [voxhammer, nano3d], stylization [stylesculptor], and scene modeling [hiscene]. A naive strategy for 3D morphing with SLAT involves first generating a 2D morphing sequence and then lifting each frame independently into 3D using Trellis. However, as illustrated in Fig. 2-(b), this approach fails to ensure temporal consistency due to frame-wise generation and often misrepresents complex 3D deformations. A strategy that is more deeply integrated with the generative framework is to directly interpolate the initial noise and conditional features of Trellis, similar to generative-prior-based morphing methods [diffmorpher, 3Dmorpher]. Yet, as shown in Fig. 2-(c), this strategy lacks explicit constraints on structural plausibility and temporal continuity, often producing suboptimal morphing quality. In summary, while the SLAT representation offers compelling opportunities for 3D morphing, achieving smooth, high-fidelity, and temporally coherent 3D morphing within modern SLAT-based generative frameworks remains an open and pressing challenge.
@@ 2 Related Work
2D Morphing. Image morphing [2Dmorphing1, 2Dmorphing2, 2Dmorphing3, 2Dmorphing4, 2D3Dsurvey] has long been studied in computer vision and graphics due to its broad applications. Traditional methods [2Dmorphing1, 2Dmorphing2, 2Dmorphing3, 2Dmorphing4] rely on hand-crafted features to establish image correspondences, followed by gradual blending to produce smooth transitions. However, they often fail to generate plausible, novel content and are prone to artifacts such as ghosting. With the advent of deep learning, data-driven approaches [2Ddatamorphing1, 2Ddatamorphing2] have improved results by learning 2D generative priors. Nevertheless, their limited model capacity makes cross-category morphing challenging. The recent success of diffusion models [diffusion_ddim, diffusion_ddpm, diffusion_sde] has enabled high-fidelity and diverse image generation. By leveraging pre-trained text-to-image models [stable_diffusion], several works [interpolating, Impus, aid, diffmorpher, freemorph] have significantly enhanced morphing quality across different categories. Despite these advances, extending such smooth and semantically coherent transformations to 3D content remains an open challenge.
@@
3D Morphing. Most 3D morphing methods rely on correspondences between source and target shapes, followed by the interpolation of matched 3D primitives. The core difficulty lies in accurate 3D matching [3Dmatch_survey1, 3Dmatch_survey2, 3Dmatch_survey3, 3Dmatch_survey4]. Early methods approach this from an axiomatic perspective, incorporating theories such as optimal transport [opt_trans1, opt_trans2, morphflow] and functional maps [functionmap1, functionmap2, functionmap3, functionmap4] to compute correspondences within the same category. In the deep learning era, one line of work [datamap1, datamap2, densematcher] adopts a data-driven paradigm, enabling the prediction of 3D correspondences using curated 3D datasets. Another category [featuremap1, featuremap3, featuremap4, featuremap5, gaussianmorphing] leverages powerful 2D feature extractors [clip, dinov2, stable_diffusion] to establish 3D correspondences in a zero-shot manner. However, these methods struggle to reliably find correspondences across diverse object categories, limiting the practicality of matching-based 3D morphing.
@@
3DMorpher [3Dmorpher] improves cross-category morphing performance by incorporating 3DGS-based 3D generative priors. However, it cannot handle complex geometries due to limitations in its base generator. Moreover, 3DGS-based outputs are not compatible with most commercial 3D software [blender, maya, unrealengine]. In this work, inspired by Trellis [trellis]—a milestone in 3D generation—we propose a series of simple yet effective modules based on its SLAT representation to achieve aesthetically pleasing and semantically coherent 3D morphing without any training.
@@
3D Generative Models. Advances in 2D diffusion models [diffusion_ddim, diffusion_ddpm, diffusion_sde, dit, sit, rectified_flow] and the availability of large-scale 3D datasets [objaverse, objaverse_xl] have driven rapid progress in 3D generation [lrm, syncdreamer, wonder3d, crm, instantmesh, lgm, unique3d, gaussiancube, 3DShape2VecSet, clay, michelangelo, triposr, triposg, triposf, sparc3d, ultra3d, direct3d, direct3d_s2, 3dtopia_Xl, hunyuan3d_v1, hunyuan3d_v2, hunyuan3d_v2.5, hunyuan3d_omni, dora, craftsman3d, gaussiananything, trellis, hi3dgen] Among these, Trellis [trellis] stands out as a groundbreaking method for native 3D generation. Its proposed Structured LATent (SLAT) representation not only efficiently encodes rich visual features but is also easily modeled and understood within modern generative frameworks. Crucially, SLAT possesses a regular and explicit structure, offering superior extensibility compared to implicit or irregular representations such as VecSet [3DShape2VecSet], NeRF [nerf], or 3DGS [3dgs]. A growing body of follow-up work demonstrates that SLAT can be directly transferred—without retraining—to various downstream tasks, including 3D editing [voxhammer, nano3d], stylization [stylesculptor], correspondence estimation [HNSR], articulated object modeling [FreeArt3D], scene generation [hiscene], and world synthesis [syncity, TRELLISWorld]. Despite its strong generalization ability, SLAT has not yet been explored for 3D morphing. In this paper, we bridge this gap by analyzing fusion rules of SLAT features in cross-/self-attention blocks, and propose lightweight yet solid modules and strategies that fully unlock the potential of SLAT for high-quality, training-free 3D morphing.
@@ 3.1 Preliminaries
Trellis [trellis] is a robust feed-forward 3D generative model that produces diverse, high-fidelity 3D assets from text or image prompts. Its core encodes 3D models into the Structured LATent (SLAT) representation, i.e., a set of local latent vectors $\{(z_{i},p_{i})\}_{i=1}^{L}$ anchored at active sparse voxels $p_{i}$ on the object surface, where each $z_{i}\in\mathbb{R}^{C}$ captures fine-scale geometry and appearance. Trellis employs a two-stage generation pipeline based on rectified flow models [rectified_flow] tailored for SLAT generation: (1) Sparse Structure (SS) Stage: SS flow transformer estimates a $64^{3}$ voxel grid to identify the sparse structure $P=\{p_{i}\}_{i=1}^{L}$ representing the global shape of assets. (2) Structured Latent (SLAT) Stage: SLAT flow transformer predicts the local latent vectors $Z=\{z_{i}\}_{i=1}^{L}$ for the voxels identified in the SS stage, producing rich geometry and texture details. Finally, the generated SLAT is decoded into standard 3D representations such as meshes, NeRF [nerf], or 3DGS [3dgs]. In this work, we adopt the image-to-3D variant of Trellis [trellis], as image conditions facilitate modeling finer details. Notably, as shown in Sec. 4.5, our method generalizes seamlessly to other SLAT-based models with different input modalities.
Attention [transformer] mechanism is a fundamental component in modern generative models [dit, sit] and is formulated as:
@@ 3.2 Overview
Given a source object $x^{\text{src}}$ and a target object $x^{\text{tgt}}$, our goal is to generate a smooth, high-quality morphing sequence $\{x^{n}\}_{n=0}^{N}$ that gradually transforms $x^{\text{src}}$ into $x^{\text{tgt}}$ using the SLAT representation from Trellis [trellis]. Both objects may be real 3D assets or Trellis-synthesized. For real assets, we obtain their initial noised latents $f^{\text{src}}_{\text{init}}$ and $f^{\text{tgt}}_{\text{init}}$, and image conditions $c^{\text{src}}$, $c^{\text{tgt}}$ via 3D inversion [voxhammer]. Here, the latent $f=(f_{ss},f_{slat})$ for each object corresponds to Trellis’s SS and SLAT stages. For Trellis-generated assets, we reuse cached initial features and conditions from the original generation. The initial noisy feature for frame $n$, $f^{n}_{\text{init}}$, is computed by spherical interpolation [diffmorpher] with a deformation weight $\alpha^{n}\in[0,1]$, ensuring $x^{0}=x^{\text{src}}(\alpha^{0}=0)$ and $x^{N}=x^{\text{tgt}}(\alpha^{N}=1)$. We set $N=49$ (50 frames) with linearly spaced $\alpha^{n}=n/N$. Further details are provided in our Supp. Mat. Fig. 3-(a) shows the MorphAny3D overview. Based on observed SLAT feature blending patterns in cross-/self-attention blocks (Sec. 3.3, Fig. 4), we introduce two key components to improve morphing plausibility and temporal coherence: Morphing Cross-Attention (Sec. 3.4, Fig. 3-(b)) and Temporal-Fused Self-Attention (Sec. 3.5, Fig. 3-(c)). We also propose an orientation correction strategy (Sec. 3.6, Fig. 3-(d)) derived from statistical analysis of orientation distributions in Trellis outputs to suppress abrupt pose shifts.
@@ 3.3 SLAT Fusion Patterns in Attention Mechanism
As shown in Fig. 2, naively applying SLAT to 3D morphing yields poor transitions, highlighting the need to understand how SLAT features influence the morphing process. Prior work [diffmorpher, 3Dmorpher] shows that fusing source and target keys and values in attention greatly improves 2D/3D morphing quality and continuity:
@@ 3.4 Morphing Cross-Attention (MCA)
We hypothesize that the local irrational structure in KV-Fused CA is caused by patch-wise blending of source and target 2D semantics, which introduces ambiguous conditions that mislead the generation model. Specifically, the blended keys and values in cross-attention are derived from patch-wise DINOv2 [dinov2] features. However, spatially aligned patches between source and target images often lack semantic correspondence, leading to distorted outputs. To verify this hypothesis, we visualize the cross-attention maps of head SLAT features (marked by red stars) in Fig. 5. The attention maps from vanilla CA (first two columns) correctly focus on the head regions in the input conditions (indicated by pink stars). In contrast, KV-Fused CA (third column) erroneously attends to background regions (highlighted by an orange box), subsequently using semantically mismatched features to guide head SLAT generation—resulting in local structural distortions. To address this issue, we propose Morphing Cross-Attention (MCA) (Fig. 3-(b)). Instead of blending keys and values prior to the attention computation, MCA computes separate attention outputs for the source and target objects and then combines them to produce the final feature $f^{n}_{MCA}$:
@@ 3.6 Orientation Correction
Morphing often suffers from abrupt orientation changes (i.e., orientation jumps)—especially in intermediate stages—causing visually jarring transitions (Fig. 6-(a)). We represent orientation via the Z-Y-X Euler angles $E=(\text{yaw},\text{pitch},\text{roll})$ (in degrees), estimated using OrientAnything [OrientAnything]. An orientation jump is flagged when the adjacent-frame angular difference $\Delta E$ exceeds 45°, a threshold accounting for estimation noise.
@@ 4.1 Experimental Settings
Implementation Details. Our method uses Image-to-3D Trellis [trellis] without retraining or hyperparameter tuning. Experiments were conducted on a single A6000 GPU. Generating each frame takes 30s with 24GB of memory.
Baselines. We compare four types: (1) Matching-based 3D morphing of 3D or SLAT features (3DInterp, SLATInterp, and use DenseMatcher [densematcher] for correspondence); (2) 2D morphing (DiffMorpher [diffmorpher], FreeMorph [freemorph]) is lifted to 3D via Trellis; (3) Direct interpolation of noise and conditions (DirectInterp); (4) Modern 3D morphing method—MorphFlow [morphflow]. Since 3DMorpher [3Dmorpher] has not released its full code, we provide only a qualitative comparison in our Supp. Mat.
Evaluation Metrics. We evaluate 50 diverse source-target pairs from real 3D datasets [voxhammer, gso] and Trellis-generated assets using four metrics: (a) FID [fid] measures visual plausibility by comparing rendered frames to originals; (b) Perceptual Path Length (PPL) and Perceptual Distance Variance (PDV) [diffmorpher] assess transition smoothness and temporal homogeneity via average perceptual change and its variance between adjacent frames; (c) Aesthetics Scores (AS) uses vision-language models [chatgpt, gemini] to rank visual appeal; (d) User Preference (UP) is gathered through a user study evaluating quality, smoothness, and realism.

### v2

@@ 1 Introduction
Morphing [76, 77, 43] refers to the visual effect of seamlessly transforming a source into a target through a smooth, plausible, and aesthetic deformation sequence. This foundational technique is widely used in animation, film, and game design to enhance creative expression. Based on input dimensionality, morphing is categorized into 2D morphing [36, 2, 90, 4] and 3D morphing [17, 19, 70, 38, 83]. While 2D morphing has advanced significantly with diffusion models [64, 30, 65, 59], 3D morphing remains challenging due to the inherent complexity of modeling smooth and reasonable deformations in three dimensions.
@@
Most existing 3D morphing approaches follow a two-stage paradigm: (1) establishing dense correspondences [71, 67] between source and target objects using hand-crafted sparse landmarks [17, 60], functional maps [57], optimal transport [63], data priors [19], or extracted 2D features [50]; and (2) interpolating between these correspondences to generate intermediate shapes. While effective in constrained settings, such matching-based methods face several critical limitations. First, they typically prioritize geometric deformation while ignoring the concurrent evolution of textures. Second, they often exhibit limited generalization, particularly in cross-category transformations (e.g., morphing a chair into a car). As shown in Fig. 2-(a), inaccurate correspondence estimation frequently leads to structurally implausible morphing results.
@@
Recent advances in feed-forward 3D generation frameworks [68, 47, 88, 92, 91, 81, 79] have substantially improved the quality and fidelity of 3D synthesis. Notably, Trellis [81] achieves diverse and high-fidelity 3D generation through its Structured Latent (SLAT) representation. Owing to SLAT’s explicit and regular structure, it shows strong potential for training-free applications such as 3D editing [37, 87], stylization [55], and scene modeling [13]. A naive strategy for 3D morphing with SLAT involves first generating a 2D morphing sequence and then lifting each frame independently into 3D using Trellis. However, as illustrated in Fig. 2-(b), this approach fails to ensure temporal consistency due to frame-wise generation and often misrepresents complex 3D deformations. A strategy that is more deeply integrated with the generative framework is to directly interpolate the initial noise and conditional features of Trellis, similar to generative-prior-based morphing methods [90, 83]. Yet, as shown in Fig. 2-(c), this strategy lacks explicit constraints on structural plausibility and temporal continuity, often producing suboptimal morphing quality. In summary, while the SLAT representation offers compelling opportunities for 3D morphing, achieving smooth, high-fidelity, and temporally coherent 3D morphing within modern SLAT-based generative frameworks remains an open and pressing challenge.
@@ 2 Related Work
2D Morphing. Image morphing [36, 95, 62, 42, 43] has long been studied in computer vision and graphics due to its broad applications. Traditional methods [36, 95, 62, 42] rely on hand-crafted features to establish image correspondences, followed by gradual blending to produce smooth transitions. However, they often fail to generate plausible, novel content and are prone to artifacts such as ghosting. With the advent of deep learning, data-driven approaches [2, 23] have improved results by learning 2D generative priors. Nevertheless, their limited model capacity makes cross-category morphing challenging. The recent success of diffusion models [64, 30, 65] has enabled high-fidelity and diverse image generation. By leveraging pre-trained text-to-image models [59], several works [73, 85, 27, 90, 4] have significantly enhanced morphing quality across different categories. Despite these advances, extending such smooth and semantically coherent transformations to 3D content remains an open challenge.
@@
3D Morphing. Most 3D morphing methods rely on correspondences between source and target shapes, followed by the interpolation of matched 3D primitives. The core difficulty lies in accurate 3D matching [71, 67, 61, 12]. Early methods approach this from an axiomatic perspective, incorporating theories such as optimal transport [63, 20, 70] and functional maps [53, 58, 57, 18] to compute correspondences within the same category. In the deep learning era, one line of work [19, 3, 94] adopts a data-driven paradigm, enabling the prediction of 3D correspondences using curated 3D datasets. Another category [25, 16, 66, 44, 38] leverages powerful 2D feature extractors [56, 52, 59] to establish 3D correspondences in a zero-shot manner. However, these methods struggle to reliably find correspondences across diverse object categories, limiting the practicality of matching-based 3D morphing.
@@
3DMorpher [83] improves cross-category morphing performance by incorporating 3DGS-based 3D generative priors. However, it cannot handle complex geometries due to limitations in its base generator. Moreover, 3DGS-based outputs are not compatible with most commercial 3D software [24, 1, 22]. In this work, inspired by Trellis [81]—a milestone in 3D generation—we propose a series of simple yet effective modules based on its SLAT representation to achieve aesthetically pleasing and semantically coherent 3D morphing without any training.
@@
3D Generative Models. Advances in 2D diffusion models [64, 30, 65, 54, 48, 45] and the availability of large-scale 3D datasets [11, 10] have driven rapid progress in 3D generation [31, 46, 47, 75, 82, 68, 78, 89, 88, 91, 93, 69, 40, 28, 41, 8, 79, 80, 9, 84, 92, 34, 32, 7, 39, 35, 81, 86] Among these, Trellis [81] stands out as a groundbreaking method for native 3D generation. Its proposed Structured LATent (SLAT) representation not only efficiently encodes rich visual features but is also easily modeled and understood within modern generative frameworks. Crucially, SLAT possesses a regular and explicit structure, offering superior extensibility compared to implicit or irregular representations such as VecSet [88], NeRF [49], or 3DGS [33]. A growing body of follow-up work demonstrates that SLAT can be directly transferred—without retraining—to various downstream tasks, including 3D editing [37, 87], stylization [55], correspondence estimation [15], articulated object modeling [5], scene generation [13], and world synthesis [21, 6]. Despite its strong generalization ability, SLAT has not yet been explored for 3D morphing. In this paper, we bridge this gap by analyzing fusion rules of SLAT features in cross-/self-attention blocks, and propose lightweight yet solid modules and strategies that fully unlock the potential of SLAT for high-quality, training-free 3D morphing.
@@ 3.1 Preliminaries
Trellis [81] is a robust feed-forward 3D generative model that produces diverse, high-fidelity 3D assets from text or image prompts. Its core encodes 3D models into the Structured LATent (SLAT) representation, i.e., a set of local latent vectors $\{(z_{i},p_{i})\}_{i=1}^{L}$ anchored at active sparse voxels $p_{i}$ on the object surface, where each $z_{i}\in\mathbb{R}^{C}$ captures fine-scale geometry and appearance. Trellis employs a two-stage generation pipeline based on rectified flow models [45] tailored for SLAT generation: (1) Sparse Structure (SS) Stage: SS flow transformer estimates a $64^{3}$ voxel grid to identify the sparse structure $P=\{p_{i}\}_{i=1}^{L}$ representing the global shape of assets. (2) Structured Latent (SLAT) Stage: SLAT flow transformer predicts the local latent vectors $Z=\{z_{i}\}_{i=1}^{L}$ for the voxels identified in the SS stage, producing rich geometry and texture details. Finally, the generated SLAT is decoded into standard 3D representations such as meshes, NeRF [49], or 3DGS [33]. In this work, we adopt the image-to-3D variant of Trellis [81], as image conditions facilitate modeling finer details. Notably, as shown in Sec. 4.5, our method generalizes seamlessly to other SLAT-based models with different input modalities.
Attention [72] mechanism is a fundamental component in modern generative models [54, 48] and is formulated as:
@@ 3.2 Overview
Given a source object $x^{\text{src}}$ and a target object $x^{\text{tgt}}$, our goal is to generate a smooth, high-quality morphing sequence $\{x^{n}\}_{n=0}^{N}$ that gradually transforms $x^{\text{src}}$ into $x^{\text{tgt}}$ using the SLAT representation from Trellis [81]. Both objects may be real 3D assets or Trellis-synthesized. For real assets, we obtain their initial noised latents $f^{\text{src}}_{\text{init}}$ and $f^{\text{tgt}}_{\text{init}}$, and image conditions $c^{\text{src}}$, $c^{\text{tgt}}$ via 3D inversion [37]. Here, the latent $f=(f_{ss},f_{slat})$ for each object corresponds to Trellis’s SS and SLAT stages. For Trellis-generated assets, we reuse cached initial features and conditions from the original generation. The initial noisy feature for frame $n$, $f^{n}_{\text{init}}$, is computed by spherical interpolation [90] with a deformation weight $\alpha^{n}\in[0,1]$, ensuring $x^{0}=x^{\text{src}}(\alpha^{0}=0)$ and $x^{N}=x^{\text{tgt}}(\alpha^{N}=1)$. We set $N=49$ (50 frames) with linearly spaced $\alpha^{n}=n/N$. Further details are provided in our Supp. Mat. Fig. 3-(a) shows the MorphAny3D overview. Based on observed SLAT feature blending patterns in cross-/self-attention blocks (Sec. 3.3, Fig. 4), we introduce two key components to improve morphing plausibility and temporal coherence: Morphing Cross-Attention (Sec. 3.4, Fig. 3-(b)) and Temporal-Fused Self-Attention (Sec. 3.5, Fig. 3-(c)). We also propose an orientation correction strategy (Sec. 3.6, Fig. 3-(d)) derived from statistical analysis of orientation distributions in Trellis outputs to suppress abrupt pose shifts.
@@ 3.3 SLAT Fusion Patterns in Attention Mechanism
As shown in Fig. 2, naively applying SLAT to 3D morphing yields poor transitions, highlighting the need to understand how SLAT features influence the morphing process. Prior work [90, 83] shows that fusing source and target keys and values in attention greatly improves 2D/3D morphing quality and continuity:
@@ 3.4 Morphing Cross-Attention (MCA)
We hypothesize that the local irrational structure in KV-Fused CA is caused by patch-wise blending of source and target 2D semantics, which introduces ambiguous conditions that mislead the generation model. Specifically, the blended keys and values in cross-attention are derived from patch-wise DINOv2 [52] features. However, spatially aligned patches between source and target images often lack semantic correspondence, leading to distorted outputs. To verify this hypothesis, we visualize the cross-attention maps of head SLAT features (marked by red stars) in Fig. 5. The attention maps from vanilla CA (first two columns) correctly focus on the head regions in the input conditions (indicated by pink stars). In contrast, KV-Fused CA (third column) erroneously attends to background regions (highlighted by an orange box), subsequently using semantically mismatched features to guide head SLAT generation—resulting in local structural distortions. To address this issue, we propose Morphing Cross-Attention (MCA) (Fig. 3-(b)). Instead of blending keys and values prior to the attention computation, MCA computes separate attention outputs for the source and target objects and then combines them to produce the final feature $f^{n}_{MCA}$:
@@ 3.6 Orientation Correction
Morphing often suffers from abrupt orientation changes (i.e., orientation jumps)—especially in intermediate stages—causing visually jarring transitions (Fig. 6-(a)). We represent orientation via the Z-Y-X Euler angles $E=(\text{yaw},\text{pitch},\text{roll})$ (in degrees), estimated using OrientAnything [74]. An orientation jump is flagged when the adjacent-frame angular difference $\Delta E$ exceeds 45°, a threshold accounting for estimation noise.
@@ 4.1 Experimental Settings
Implementation Details. Our method uses Image-to-3D Trellis [81] without retraining or hyperparameter tuning. Experiments were conducted on a single A6000 GPU. Generating each frame takes 30s with 24GB of memory.
Baselines. We compare four types: (1) Matching-based 3D morphing of 3D or SLAT features (3DInterp, SLATInterp, and use DenseMatcher [94] for correspondence); (2) 2D morphing (DiffMorpher [90], FreeMorph [4]) is lifted to 3D via Trellis; (3) Direct interpolation of noise and conditions (DirectInterp); (4) Modern 3D morphing method—MorphFlow [70]. Since 3DMorpher [83] has not released its full code, we provide only a qualitative comparison in our Supp. Mat.
Evaluation Metrics. We evaluate 50 diverse source-target pairs from real 3D datasets [37, 14] and Trellis-generated assets using four metrics: (a) FID [29] measures visual plausibility by comparing rendered frames to originals; (b) Perceptual Path Length (PPL) and Perceptual Distance Variance (PDV) [90] assess transition smoothness and temporal homogeneity via average perceptual change and its variance between adjacent frames; (c) Aesthetics Scores (AS) uses vision-language models [51, 26] to rank visual appeal; (d) User Preference (UP) is gathered through a user study evaluating quality, smoothness, and realism.